OpenStudy (anonymous):

5. Which of the following side lengths do NOT form a right triangle? A. 5 cm, 12 cm and 13 cm B. 8 cm, 15 cm and 17 cm C. 9 in, 40 in and 41 in D. 20 in, 21 in and 28 in 6.To get from your house to your grandparent’s house you must avoid walking through a pond. To avoid the pond, you walk 0.5 miles south and 1.2 miles east. How many miles would be saved if it were possible to walk through the pond? A. 0.4 miles B. 1.69 miles C. 3.0 miles D. 6.4 miles

OpenStudy (anonymous):

To start you just need to use Pythagorean Theorem. a^2+b^2=c^2

OpenStudy (mathlegend):

@Ambbiiee what answer did you come up with for 5?

OpenStudy (anonymous):

The answer to number 5 is B

OpenStudy (mathlegend):

@BronxWarrior lets not just give answers. @Ambbiiee has to do the work. It would be best if Ambbiiee actually showed us the process.

OpenStudy (anonymous):

i picked B but my teacher said it was wrong @MathLegend

OpenStudy (mathlegend):

Remember it says which do NOT form a right triangle. If \[a ^{2} + b ^{2} = c ^{2}\] Then it is a Right Triangle. You are looking for one where the operations do not add up.

OpenStudy (anonymous):

wait no

OpenStudy (mathlegend):

@Marielyne don't give out answers.

OpenStudy (mathlegend):

Especially wrong ones.

OpenStudy (anonymous):

is it D?

OpenStudy (mathlegend):

Yes!

OpenStudy (mathlegend):

400 + 441 "does not equal" = 784

OpenStudy (mathlegend):

So, great job!

OpenStudy (anonymous):

so 5 is D

OpenStudy (anonymous):

can you help me with 6

OpenStudy (mathlegend):

Yes and for 6 we still have to use the Pythagorean Theorem.

OpenStudy (mathlegend):

Lets try to draw a picture based off of what it says... here are two key phrases... "you walk 0.5 miles south"

OpenStudy (mathlegend):

So how would you draw a line going south? Can you draw that for me? Would it be vertical or horizontal?

OpenStudy (anonymous):

horizontal

OpenStudy (mathlegend):



OpenStudy (mathlegend):

@Jhannybean ... I was trying to get @Ambbiiee to see how I was going to make the picture. Thanks for jumping to conclusions.

OpenStudy (mathlegend):

@Ambbiiee needs to be the one to be able to interpret the words.

OpenStudy (mathlegend):

@Jhannybean go ahead and finish the problem. @Ambbiiee I am sorry but @Jhannybean will finish helping you.

OpenStudy (mathlegend):

No go ahead.

OpenStudy (whpalmer4):

why do you keep reposting this question?

OpenStudy (mathlegend):

@Jhannybean please come and finish the problem.